# AgenticSciML: Collaborative Multi-Agent Systems for Emergent Discovery in Scientific Machine Learning

### v1

@@ B.2 List of Knowledge Base Entries
1 Adaptive Act. Adaptive activation functions for Burgers equation [jagtap2020adaptive]
2 Local Adaptive Act. Locally adaptive activation with slope recovery [jagtap2020locally]
3 DeepONet Antiderivative operator learning [lu2021learning, toscano2024deeponet]
4 PI-DeepONet Physics-informed antiderivative operator learning [wang2021learning, toscano2024deeponet]
5 PI-DeepONet Diffusion-reaction systems [wang2021learning, toscano2024deeponet]
6 MoE-PINN Burgers equation with mixture-of-experts [bischof2022mixture]
7 MoE-PINN Poisson equation on L-shaped domain [bischof2022mixture]
8 Self-Adaptive PINN Allen-Cahn equation [mcclenny2023self]
9 Self-Adaptive PINN Burgers equation [mcclenny2023self]
10 Self-Adaptive PINN Helmholtz equation [mcclenny2023self]
11 PINN Burgers equation forward problem [raissi2019physics]
12 PINN Burgers equation inverse problem [raissi2019physics]
13 PINN Diffusion equation with source term [raissi2019physics]
14 PINN Simple ordinary differential equation [raissi2019physics]
15 PINN 1D Poisson equation with Dirichlet BC [raissi2019physics]
16 hp-VPINN Advection-diffusion inverse problem [kharazmi2021hp]
17 hp-VPINN 1D Poisson equation [kharazmi2021hp]
18 hp-VPINN 2D Poisson equation [kharazmi2021hp]
19 hp-VPINN 2D Poisson equation on L-shaped domain [kharazmi2021hp]
20 gPINN Diffusion-reaction equation [yu2022gradient]
21 Delta-PINN Complex geometries [costabal2024delta]
22 gPINN Allen-Cahn equation with adaptive refinement [yu2022gradient]
23 gPINN Burgers’ equation [yu2022gradient]
24 gPINN Inverse diffusion-reaction system [yu2022gradient]
25 gPINN Function approximation [yu2022gradient]
26 gPINN 1D Poisson equation [yu2022gradient]
27 gPINN 2D Poisson equation [yu2022gradient]
28 NSFnet Incompressible Navier-Stokes equations [jin2021nsfnets]
29 ev-NSFnet Entropy-viscosity regularized NSFnet [jin2021nsfnets]
30 hPINN Inverse design with hard constraints [lu2021physics]
31 hPINN Stokes flow topology optimization [lu2021physics]
32 DeepONet 2D advection equations [lu2022comprehensive, lu2021learning]
33 FNO 2D advection equations [lu2022comprehensive, li2020fourier]
34 DeepONet 1D Burgers’ equation [lu2022comprehensive, lu2021learning]
35 FNO 1D Burgers’ equation [lu2022comprehensive, li2020fourier]
36 DeepONet 2D Darcy flow with piecewise constant permeability [lu2022comprehensive, lu2021learning]
37 FNO 2D Darcy flow [lu2022comprehensive, li2020fourier]
38 DeepONet 2D Darcy flow on triangular domain with notch [lu2022comprehensive, lu2021learning]
39 FNO Irregular geometries [lu2022comprehensive, li2020fourier]
40 POD-DeepONet Instability wave propagation [lu2022comprehensive, lu2021learning]
41 FNO Wave propagation with non-uniform grids [lu2022comprehensive, li2020fourier]
42 ICON In-context operator networks [yang2023context]
43 U-FNO Gas saturation in CO2 storage [wen2022u]
44 U-FNO Pressure buildup in CO2 storage [wen2022u]
45 CNO Allen-Cahn equation [raonic2023convolutional]
46 Fourier-DeepONet Full waveform inversion (CurveFault-A) [zhu2023fourier]
47 Fourier-DeepONet Full waveform inversion (CurveVel-A) [zhu2023fourier]
48 Fourier-DeepONet Full waveform inversion (FlatVel-B) [zhu2023fourier]
49 Fourier-DeepONet Full waveform inversion (Style-A) [zhu2023fourier]
50 CNO Smooth transport equation [raonic2023convolutional]
51 CNO Darcy flow [raonic2023convolutional]
52 CNO Discontinuous transport equation [raonic2023convolutional]
53 CNO Compressible Euler equations (airfoil) [raonic2023convolutional]
54 CNO Navier-Stokes shear layer [raonic2023convolutional]
55 CNO Wave equation [raonic2023convolutional]
56 CNO-FM Foundation model for multiple PDEs [raonic2023convolutional]
57 Oformer Compressible flow around airfoils [li2022transformer]
58 Oformer Burgers’ equation [li2022transformer]
59 Oformer Darcy flow [li2022transformer]
60 Oformer Electrostatics Poisson equation [li2022transformer]
61 Oformer Magnetostatics Poisson equation [li2022transformer]
62 Oformer 2D Navier-Stokes equations [li2022transformer]
63 GNOT 2D Darcy flow [hao2023gnot]
64 GNOT 2D elasticity [hao2023gnot]
65 GNOT 2D multi-scale heat conduction [hao2023gnot]
66 GNOT 3D multi-physics heatsink [hao2023gnot]
67 GNOT 2D electromagnetic inductor [hao2023gnot]
68 GNOT 2D transonic airfoil flow [hao2023gnot]
69 GNOT 2D Navier-Stokes steady-state [hao2023gnot]
70 GNOT 2D time-dependent Navier-Stokes [hao2023gnot]
Appendix C Agent Prompts
@@
C.1 Data analyst agent
@@ C.7 Debugger agent
C.8 Result analyst agent

### v2

@@ 4 Summary and Discussion
1 Adaptive Act. Adaptive activation functions for Burgers equation [15]
2 Local Adaptive Act. Locally adaptive activation with slope recovery [14]
3 DeepONet Antiderivative operator learning [27, 44]
4 PI-DeepONet Physics-informed antiderivative operator learning [46, 44]
5 PI-DeepONet Diffusion-reaction systems [46, 44]
6 MoE-PINN Burgers equation with mixture-of-experts [2]
7 MoE-PINN Poisson equation on L-shaped domain [2]
8 Self-Adaptive PINN Allen-Cahn equation [31]
9 Self-Adaptive PINN Burgers equation [31]
10 Self-Adaptive PINN Helmholtz equation [31]
11 PINN Burgers equation forward problem [37]
12 PINN Burgers equation inverse problem [37]
13 PINN Diffusion equation with source term [37]
14 PINN Simple ordinary differential equation [37]
15 PINN 1D Poisson equation with Dirichlet BC [37]
16 hp-VPINN Advection-diffusion inverse problem [20]
17 hp-VPINN 1D Poisson equation [20]
18 hp-VPINN 2D Poisson equation [20]
19 hp-VPINN 2D Poisson equation on L-shaped domain [20]
20 gPINN Diffusion-reaction equation [51]
21 Delta-PINN Complex geometries [6]
22 gPINN Allen-Cahn equation with adaptive refinement [51]
23 gPINN Burgers’ equation [51]
24 gPINN Inverse diffusion-reaction system [51]
25 gPINN Function approximation [51]
26 gPINN 1D Poisson equation [51]
27 gPINN 2D Poisson equation [51]
28 NSFnet Incompressible Navier-Stokes equations [19]
29 ev-NSFnet Entropy-viscosity regularized NSFnet [19]
30 hPINN Inverse design with hard constraints [29]
31 hPINN Stokes flow topology optimization [29]
32 DeepONet 2D advection equations [28, 27]
33 FNO 2D advection equations [28, 23]
34 DeepONet 1D Burgers’ equation [28, 27]
35 FNO 1D Burgers’ equation [28, 23]
36 DeepONet 2D Darcy flow with piecewise constant permeability [28, 27]
37 FNO 2D Darcy flow [28, 23]
38 DeepONet 2D Darcy flow on triangular domain with notch [28, 27]
39 FNO Irregular geometries [28, 23]
40 POD-DeepONet Instability wave propagation [28, 27]
41 FNO Wave propagation with non-uniform grids [28, 23]
42 ICON In-context operator networks [50]
43 U-FNO Gas saturation in CO2 storage [48]
44 U-FNO Pressure buildup in CO2 storage [48]
45 CNO Allen-Cahn equation [38]
46 Fourier-DeepONet Full waveform inversion (CurveFault-A) [56]
47 Fourier-DeepONet Full waveform inversion (CurveVel-A) [56]
48 Fourier-DeepONet Full waveform inversion (FlatVel-B) [56]
49 Fourier-DeepONet Full waveform inversion (Style-A) [56]
50 CNO Smooth transport equation [38]
51 CNO Darcy flow [38]
52 CNO Discontinuous transport equation [38]
53 CNO Compressible Euler equations (airfoil) [38]
54 CNO Navier-Stokes shear layer [38]
55 CNO Wave equation [38]
56 CNO-FM Foundation model for multiple PDEs [38]
57 Oformer Compressible flow around airfoils [22]
58 Oformer Burgers’ equation [22]
59 Oformer Darcy flow [22]
60 Oformer Electrostatics Poisson equation [22]
61 Oformer Magnetostatics Poisson equation [22]
62 Oformer 2D Navier-Stokes equations [22]
63 GNOT 2D Darcy flow [12]
64 GNOT 2D elasticity [12]
65 GNOT 2D multi-scale heat conduction [12]
66 GNOT 3D multi-physics heatsink [12]
67 GNOT 2D electromagnetic inductor [12]
68 GNOT 2D transonic airfoil flow [12]
69 GNOT 2D Navier-Stokes steady-state [12]
70 GNOT 2D time-dependent Navier-Stokes [12]
S3 Agent Prompts
@@ 70 GNOT 2D time-dependent Navier-Stokes [12]
S3.1 Data analyst agent
@@
S3.8 Result analyst agent